OpenStudy (anonymous):

You and your friend are saving for college. You have $50 and are adding $10 each week to your savings. Your friend has $20 and he is adding $20 each week to his savings. a. What system of equations would be a good model for this situation? Let x be number of weeks and y be the number of dollars saved. b. Graph your system of equations. c. Use your graph to determine when you and your friend will have the same amount of money saved. Explain your answer.

Directrix (directrix):

Do you think this is a 2 equations in 2 unknowns situation or 1 equation in 1 unknown?

OpenStudy (anonymous):

No idea

OpenStudy (anonymous):

Okay thank you

Directrix (directrix):

You (in the problem) $50 + 10*x where x is the numer of weeks money is saved Friend: $20 + 20*x

Directrix (directrix):

YOU y = $50 + 10*x FRIEND y = $20 + 20*x

Directrix (directrix):

@Juliasawyer Are we to show the graph or just give the results of the graph?

OpenStudy (anonymous):

Show the graph!

Directrix (directrix):

Okay. Just a sec.

Directrix (directrix):

Time to analyze this. See attachment.

OpenStudy (anonymous):

That makes no sense

Directrix (directrix):

@Juliasawyer What is the value of x

Directrix (directrix):

On the graph, the point of intersection gives the number of weeks and the amount both will have saved (same amount)

Directrix (directrix):

Hey, you said this problem is important so let's knock it out.

Directrix (directrix):

x = ?

Directrix (directrix):

@Juliasawyer

OpenStudy (anonymous):

(3,80)

Directrix (directrix):

Correct. After 3 weeks, each will have $80

Directrix (directrix):

Did you get the graph?

OpenStudy (anonymous):

Where else should each line hit the x and y axis on the graph?

Directrix (directrix):

If x = 0, then y = $50 + 10*x crosses the y-axis at 50. 50 + 10*0 = 50

Directrix (directrix):

If y = 0, then y = $50 + 10*x crossed the x-axis at y = -5.

OpenStudy (anonymous):

Okay thank you!

Directrix (directrix):

Do you have both graphs drawn on the same set of axes?